# literature.bib

### v4

@@ 3.2. Local well-posedness
From now on, we let $\psi\in L^{2}(0,T;U)$ be arbitrary but fixed. Using Corollary 3.5, we will prove local well-posedness of the actual skeleton equation (2.8). Local well-posedness is established in Theorem 3.7. Its proof and preparatory lemma’s are analogous to [HNVWvolume3, §18.2], which was inspired by [pruss17addendum] and [pruss18JDE].
The skeleton equation does not fit in the setting of [HNVWvolume3], [pruss17addendum] or [pruss18JDE], for the reason that we only have $L^{1}$-( instead of $L^{2}$-)integrability of the term $B(\cdot,u^{\psi}(\cdot))\psi(\cdot)$ in (2.8). Besides that, our maximal regularity space $\mathrm{MR}(0,T)=C([0,T];H)\cap L^{2}(0,T;V)$ is different.
@@
Suppose that $(A,B)$ satisfies Assumption 2.2. Let $u_{0}\in H$ be fixed. Then there exist $\tilde{T},\varepsilon>0$ such that for each $v_{0}\in B_{H}(u_{0},\varepsilon)$, there exists a unique strong solution $u_{v_{0}}\in\mathrm{MR}(0,\tilde{T})$ to
@@
Theorem 3.7 will be proved using the Banach fixed point theorem, applied to the map $\Psi_{v_{0}}\colon\mathrm{MR}(0,\tilde{T})\to\mathrm{MR}(0,\tilde{T})$ defined by $\Psi_{v_{0}}(v)\coloneqq u$, where $u$ is the unique strong solution to
@@
Note that $u\in\mathrm{MR}(0,\tilde{T})$ is a strong solution to (3.10) if and only if $\Psi_{v_{0}}(u)=u$.
@@
    $\leq C_{n,T}\left(\int_{0}^{T}\|u(s)-v(s)\|_{H}^{2}\|w(s)\|_{V}^{2}\,\mathrm{d}s\right)^{\frac{1}{2}}\leq C_{n,T}\|u-v\|_{C([0,T];H)}\|w\|_{L^{2}(0,T;V)}.$
@@
In (ii)-(v), note that strong measurability of $A_{0}(\cdot,u(\cdot))w(\cdot)$, $F(\cdot,u(\cdot))$, $B_{0}(\cdot,u(\cdot))w(\cdot)$ and $G(\cdot,u(\cdot))$ holds, as was mentioned in Remark 3.6. Moreover, by symmetry in Assumption 2.2, $B_{0}$ and $G$ can be estimated in the same way as $A_{0}$ and $F$. We provide the estimates for the latter. Assumption 2.2(3) immediately yields (ii) and (iii).
@@
For (v): the following estimates can be found in the proof of [AV22variational, Prop. 4.5]. By Assumption 2.2(3) we have pointwise in $t\in[0,T]$:
@@
where $M_{\sigma,\beta,\rho}>0$ is a constant depending only on $\sigma$, $\beta$ and $\rho$ and we let $10^{0}=$ 1. In the above we used that $a\coloneqq\frac{(2\beta-1)\rho}{2-2\beta}\in[0,1]$, hence $x^{a}\leq 1+x$ for $x\geq 0$. For $j\in\{1,\ldots,m_{F}\}$, application of (3.2) gives pointwise in $t\in[0,T]$:
[…]
@@
From Lemma 3.8, we see that $\tilde{f}$ and $\tilde{g}$ defined by (3.13) lie in $L^{2}(0,\tilde{T};V^{*})$ and $L^{2}(0,\tilde{T};\mathcal{L}_{2}(U,H))$ respectively, for any $\tilde{T}>0$ and $v\in\mathrm{MR}(0,\tilde{T})$ (put $n=\|u_{0}\|_{H}\vee\|v\|_{C([0,\tilde{T}];H)}$ and apply (iii) and (v)). Thus Corollary 3.5 gives that (3.12) is well-posed, i.e. $\Psi_{v_{0}}$ is well-defined.
@@
Note that $Z_{r,\tilde{T}}(v_{0})$ is closed in $\mathrm{MR}(0,\tilde{T})$, hence complete. Eventually, we will find that $\Psi_{v_{0}}$ is contractive on some $Z_{r,\tilde{T}}(v_{0})$. Several crucial estimates will be gathered in the next lemma’s.
@@
The next lemma is analogous to [HNVWvolume3, Lem. 18.2.10].
@@
where $\alpha_{T_{1}}(\tilde{T}),\beta_{T_{1},\sigma}(\tilde{T},r)\downarrow 0$ as $\tilde{T},r\downarrow 0$ and $\alpha_{T_{1}}(\tilde{T})$ and $\beta_{T_{1},\sigma}(\tilde{T},r)$ are independent of $v_{0}$ and $v$.
@@
(3.24) $\|F(v)\|_{L^{2}(0,\tilde{T};V^{*})}+\|f\|_{L^{2}(0,\tilde{T};V^{*})}\leq\tilde{\beta}_{T_{1},\sigma}(\tilde{T},r)r+\sigma r+\gamma(\tilde{T}),$
@@
$\alpha_{T_{1}}(\tilde{T})\coloneqq 2C_{T_{1}}\alpha(\tilde{T})^{2}+\gamma(\tilde{T}),$
@@
Before we prove Theorem 3.7, we need one more lemma, a modification of [HNVWvolume3, Lemma 18.2.12].
@@
Let $u_{0}\in H$ and suppose that $(A,B)$ satisfies Assumption 2.2. For $\varepsilon_{1}$ and $r_{1}$ from Lemma 3.10, the following estimates hold for any $\tilde{T}\in(0,T_{1}]$, $\varepsilon\in(0,\varepsilon_{1}]$, $r\in(0,r_{1}]$, $v_{0},w_{0}\in B_{H}(u_{0},\varepsilon)$, $v\in Z_{r,\tilde{T}}(v_{0})$, $w\in Z_{r,\tilde{T}}(w_{0})$, $u\in\mathrm{MR}(0,\tilde{T})$ and $\sigma>0$:
@@
where $c_{T_{1}}$ is a constant and $\alpha(\tilde{T}),\beta(\tilde{T}),\gamma_{T_{1},\sigma}(\tilde{T},r)\downarrow 0$ as $\tilde{T},r\downarrow 0$. Moreover, $c_{T_{1}},\alpha(\tilde{T}),\beta(\tilde{T})$ and $\gamma_{T_{1},\sigma}(\tilde{T},r)$ are independent of $v_{0},w_{0},v$ and $w$.
@@
$\leq c_{T_{1}}\|v-w\|_{\mathrm{MR}(0,\tilde{T})}(r+\alpha(\tilde{T})),$
@@
We are now ready to prove Theorem 3.7. The proof is adapted from [HNVWvolume3, Th. 18.2.6].
@@
Let $\varepsilon_{1},r_{1}>0$ be as in Lemma 3.10 and let $\tilde{T}\in(0,T_{1}]$, $\varepsilon\in(0,\varepsilon_{1}]$, $r\in(0,r_{1}]$. As above, define $\Psi_{v_{0}}\colon\mathrm{MR}(0,\tilde{T})\to\mathrm{MR}(0,\tilde{T})$ by $\Psi_{v_{0}}(v)\coloneqq u$, where $u$ is the unique strong solution to (3.12). Recall that $u$ solves (3.10) if and only if $\Psi_{v_{0}}(u)=u$ and recall that $Z_{r,\tilde{T}}(v_{0})$ defined by (3.19) is closed in $\mathrm{MR}(0,\tilde{T})$, hence complete. We show that for $\tilde{T},\varepsilon,r$ small enough, the mapping $\Psi_{v_{0}}$ maps $Z_{r,\tilde{T}}(v_{0})$ to itself and is contractive. The Banach fixed point theorem then gives existence of a unique fixed point in $Z_{r,\tilde{T}}(v_{0})$, hence existence of a solution to (3.10). We will extend the uniqueness within $Z_{r,\tilde{T}}(v_{0})$ to uniqueness in $\mathrm{MR}(0,\tilde{T})$.
@@
with $\alpha_{T_{1}}(\tilde{T}),\beta_{T_{1},\sigma}(\tilde{T},r)\downarrow 0$ as $\tilde{T},r\downarrow 0$. Recall that $K_{T_{1}}$ from (3.8) only depends on $T_{1}$, $T$, $\|u_{0}\|_{H}$ and $\psi$, not on $v_{0}$ or $v$. Fixing first $\sigma\coloneqq(4K_{T_{1}})^{-1}$, we find
@@
It remains to show that uniqueness not only holds within $Z_{r,\tilde{T}}(v_{0})$ but also within the larger space $\mathrm{MR}(0,\tilde{T})$. Let $v,\tilde{v}\in\mathrm{MR}(0,\tilde{T})$ be strong solutions to (3.10) and suppose that $v\neq\tilde{v}$. Then we have $s\coloneqq\inf\{t\in[0,\tilde{T}]:v(t)\neq\tilde{v}\text{ in }H\}\in[0,\tilde{T})$ since $\mathrm{MR}(0,\tilde{T})=C([0,\tilde{T}];H)\cap L^{2}(0,\tilde{T};V)$ and $V\hookrightarrow H$ is injective. Moreover, $v(s)=\tilde{v}(s)\eqqcolon w_{0}$ as $v,\tilde{v}\in C([0,\tilde{T}];H)$ and $v(\cdot+s)$ and $\tilde{v}(\cdot+s)$ are strong solutions to
@@
and note that $\delta\in(0,\min\{T_{0},\tilde{T}-s\}]$ since $v(0+s)=\tilde{v}(0+s)=w_{0}=z_{w_{0}}(0)$. In particular, $\delta\in(0,T_{0}]$ and $v(\cdot+s),\tilde{v}(\cdot+s)\in Z_{r_{0},T}(w_{0})$ by definition of $\delta$. Uniqueness of solutions in $Z_{r_{0},\delta}(w_{0})$ implies that $v(\cdot+s)=\tilde{v}(\cdot+s)$ on $[0,\delta]$. Therefore $v=\tilde{v}$ on $[0,s+\delta]$, contradicting the definition of $s$. We conclude that $v=\tilde{v}$. ∎
@@
$\langle\bar{A}(t)v,v\rangle$ $\geq\langle A_{0}(t,w(t))v,v\rangle-\sigma|\!|\!|B_{0}(t,w(t))v|\!|\!|_{H}^{2}-C_{\sigma}\|\psi(t)\|_{U}^{2}\|v\|_{H}^{2}$
$\geq\theta_{n,T}\|v\|_{V}^{2}-({M}_{n,T}+C_{\sigma}\|\psi(t)\|_{U}^{2})\|v\|_{H}^{2}-\sigma C_{n,T}^{2}(1+n)^{2}\|v\|_{V}^{2}.$
Putting $\sigma\coloneqq\theta_{n,T}(2C_{n,T}^{2}(1+n)^{2})^{-1}$, the required coercivity (3.3) for $\bar{A}$ follows.
@@ 3.3. Global well-posedness
Similar to [pruss16, Chap. 5] and [HNVWvolume3, §18.2], we will extend Theorem 3.7 to a global well-posedness result by means of maximal solutions and a blow-up criterion.
@@
Note that maximal solutions are unique by definition. The proof of the next proposition is adapted from [HNVWvolume3, Th. 18.2.14, Th. 18.2.15] and [pruss16].
@@
satisfies $\bar{u}\in\mathrm{MR}(0,T_{*}+{\delta})$ and $\bar{u}$ is a strong solution to (3.10) on $[0,T_{*}+{\delta}]$, contradicting maximality of $(u_{*},T_{*})$. ∎
Using the blow-up criterion, we finally prove global well-posedness for the skeleton equation. Besides Assumption 2.2, we now also assume the coercivity condition (2.7) for the pair $(A,B)$. This condition has not been used so far, but it is also needed for the global well-posedness result for the stochastic evolution equation [AV22variational, Th. 3.5], see Theorem 2.3.
@@
$\leq\|x\|_{H}^{2}+2\int_{0}^{t}-\frac{1}{2}|\!|\!|B(s,u_{*}(s))|\!|\!|_{H}^{2}-\theta\|u_{*}(s)\|_{V}^{2}+M\|u_{*}(s)\|_{H}^{2}+|\phi(s)|^{2}$
$\qquad\qquad\qquad\qquad+|\!|\!|B(s,u_{*}(s))|\!|\!|_{H}\|\psi(s)\|_{U}\|u_{*}(s)\|_{H}\,\mathrm{d}s$
$\leq\|x\|_{H}^{2}+2\int_{0}^{t}-\frac{1}{2}|\!|\!|B(s,u_{*}(s))|\!|\!|_{H}^{2}-\theta\|u_{*}(s)\|_{V}^{2}+M\|u_{*}(s)\|_{H}^{2}+|\phi(s)|^{2}$
$\qquad\qquad\qquad\qquad+\frac{1}{2}|\!|\!|B(s,u_{*}(s))|\!|\!|_{H}^{2}+\frac{1}{2}\|\psi(s)\|_{U}^{2}\|u_{*}(s)\|_{H}^{2}\,\mathrm{d}s$
@@
where $K\colon[0,T_{*}]\to{\mathbb{R}}_{+}$ is increasing. Applying Lemma 3.8(iv) with $n=K(T_{*})<\infty$ we find that $F(u_{*})\in L^{2}(0,t;V^{*})$ for all $t\in(0,T_{*})$ and $L\coloneqq\sup_{t\in[0,T_{*})}\|F(u_{*})\|_{L^{2}(0,t;V^{*})}<\infty$. Thus, by the Monotone Convergence Theorem, $\|F(u_{*})\|_{L^{2}(0,T_{*};V^{*})}\leq L<\infty$. Similarly, $G(u_{*})\in L^{2}(0,T_{*};\mathcal{L}_{2}(U,H))$. Now we apply Corollary 3.5 with $T\coloneqq T^{*}$, $w\coloneqq u_{*}\in C([0,T_{*});H)\subset L^{\infty}(0,T;H)$ (extend by $w(T)\coloneqq u_{*}(0)$ on the Lebesgue null set $\{T\}$), $n\coloneqq K(T_{*})$ and $\bar{f}\coloneqq F(u_{*})+f\in L^{2}(0,T_{*};V^{*})$, $\bar{g}\coloneqq G(u_{*})+g\in L^{2}(0,T_{*};\mathcal{L}_{2}(U,H))$. Corollary 3.5 gives existence of a strong solution $\bar{u}\in\mathrm{MR}(0,T_{*})$ to (2.8) on $[0,T_{*}]$. By uniqueness of the maximal solution, it follows that $u_{*}|_{[0,T]}=\bar{u}|_{[0,T]}$ for all $T\in[0,T_{*})$. Hence $\lim_{t\uparrow T_{*}}u_{*}(t)=\lim_{t\uparrow T_{*}}\bar{u}(t)=\bar{u}(T_{*})\in H$, contradicting Proposition 3.15.
@@ 4.1. Weak convergence approach
    $Wf$ is normally distributed with mean zero and $\mathbb{E}[WfWg]=\langle f,g\rangle_{L^{2}({\mathbb{R}}_{+};U)}$,
@@
There exist several different definitions of a cylindrical Brownian motion or cylindrical Wiener process in the literature. Some references in our proof of the LDP use (an equivalent of) an ${\mathbb{R}}^{\infty}$-Brownian motion, defined below.
@@
An ${\mathbb{R}}^{\infty}$-Brownian motion (in $U$) is a pair $\tilde{W}\coloneqq((\beta_{k})_{k\in{\mathbb{N}}},(e_{k})_{k\in{\mathbb{N}}})$, with $(\beta_{k})_{k\in{\mathbb{N}}}$ a sequence of independent standard real-valued $(\mathcal{F}_{t})$-Brownian motions and $(e_{k})_{k\in{\mathbb{N}}}$ an orthonormal basis for $U$.
In Proposition A.4 of Appendix A, the connection between the $U$-cylindrical Brownian motion and the ${\mathbb{R}}^{\infty}$-Brownian motion is summarized, as well as their equivalent, but differently constructed stochastic integrals. The ${\mathbb{R}}^{\infty}$-Brownian motion of Definition 4.2 is e.g. used in [liurockner15], where it is called a cylindrical $Q$-Wiener process (with $Q\coloneqq I\in\mathcal{L}(U;U)$ the identity operator). Often, the notation $\tilde{W}(t)=\sum_{k\in{\mathbb{N}}}\beta_{k}(t)e_{k}$ is also used, which is only formal as the series does not converge in $L^{2}(\Omega;U)$. However, we will write $\tilde{W}=((\beta_{k})_{k\in{\mathbb{N}}},(e_{k})_{k\in{\mathbb{N}}})$.
@@
In view of the above remark, we assume that the filtration $(\mathcal{F}_{t})_{t\geq 0}$ is right-continuous and complete from now on, and we assume that $W$ is a $U$-cylindrical Brownian motion with respect to $(\mathcal{F}_{t})_{t\geq 0}$. Moreover, we fix any orthonormal basis $(e_{k})_{k\in{\mathbb{N}}}$ for $U$. We let $\tilde{W}=(\beta_{k})_{k\in{\mathbb{N}}},(e_{k})_{k\in{\mathbb{N}}})$ denote the unique ${\mathbb{R}}^{\infty}$-Brownian motion associated to $W$ from Proposition A.4, i.e. satisfying (A.6). In the upcoming proofs $\tilde{W}$ will be useful, since we will be applying the Yamada-Watanabe theorem and Girsanov’s theorem for ${\mathbb{R}}^{\infty}$-Brownian motions. Finally, from now on we fix a separable Hilbert space $U_{1}$ and a Hilbert-Schmidt inclusion $J\colon U\hookrightarrow U_{1}$. This is always possible: let $\langle u,v\rangle_{1}\coloneqq\sum_{k=1}^{\infty}\frac{1}{k}\langle u,e_{k}\rangle_{U}\langle e_{k},v\rangle_{U}$ for $u,v\in U$ and let $U_{1}\coloneqq\mathrm{completion}(U,\langle\cdot,\cdot\rangle_{1})$. We associate to $\tilde{W}$ the following $U_{1}$-valued process:
@@
By [liurockner15, Prop. 2.5.2], $\tilde{W}_{1}$ is a $Q_{1}$-Wiener process on $U_{1}$, with $Q_{1}\coloneqq JJ^{*}$. In what follows, $\tilde{W}_{1}$ denotes this $Q_{1}$-Wiener process defined by (4.2). We note that the paths of $\tilde{W}_{1}$ are in $C([0,T];U_{1})$.
@@
and for $K>0$,
[…]
The next theorem gives sufficient conditions for the LDP and is known as the weak convergence approach, which originates from [budhidupuis01, Th. 4.4]. In [matoussi21], a useful adaptation was proved. The following version is immediately derived from [matoussi21, Th. 3.2]. We will use it to prove Theorem 2.6.
Theorem 4.5.
Let $\mathcal{E}$ be a Polish space and let $(Y^{\varepsilon})_{\varepsilon>0}$ be a collection of $\mathcal{E}$-valued random variables on $(\Omega,\mathcal{F},\mathbb{P})$. Let $\tilde{W}=((\beta_{k})_{k\in{\mathbb{N}}},(e_{k})_{k\in{\mathbb{N}}})$ be an ${\mathbb{R}}^{\infty}$-Brownian motion. Let $\tilde{W}_{1}\colon\Omega\to C([0,T];U_{1})$ be the associated $Q_{1}$-Wiener process on $U_{1}$ defined by (4.2). Suppose that for $\varepsilon\geq 0$, there exist measurable maps $\mathcal{G}^{\varepsilon}\colon C([0,T];U_{1})\to\mathcal{E}$ such that
@@
    $\mathcal{G}^{0}\left(\int_{0}^{\cdot}\psi_{n}(s)\,\mathrm{d}s\right)\to\mathcal{G}^{0}\left(\int_{0}^{\cdot}\psi(s)\,\mathrm{d}s\right)\text{ in }\mathcal{E},$
@@
Conditions (ii) and (iii) imply the conditions of the original weak convergence approach of [budhidupuis01]. For the latter, instead of (ii) and (iii), one would require
@@
Here, [budhidupuis01, Th. 4.4] is applied with $Q_{1}$-Wiener process $\tilde{W}_{1}$, $H\coloneqq U_{1}$, $H_{0}\coloneqq Q_{1}^{\frac{1}{2}}(U_{1})$ and one uses that $Q_{1}^{\frac{1}{2}}(U_{1})=J(U)=U$ as a subspace of $U_{1}$, see [liurockner15, Prop. 2.5.2] (with $Q\coloneqq I$, $U_{0}\coloneqq I^{\frac{1}{2}}(U)=U$).
@@
We will verify that all conditions in Theorem 4.5 are satisfied for $Y^{\varepsilon}$ defined as the strong solution to (4.1). Condition (i) follows from the Yamada-Watanabe theorem in [rock08]. The details are given in Lemma 4.6, as well as a preparation for the proof of condition (iii).
@@
To prove the first statement, we use the Yamada-Watanabe theorem from [rock08, Th. 2.1] on $[0,T]$ with $L^{1}(0,T;V)$ replaced by $L^{2}(0,T;V)$. Let $\varepsilon>0$. For any $Y^{\varepsilon}$ with $Y^{\varepsilon}\in\mathrm{MR}(0,T)$ a.s. and for any $\xi\in L^{0}((\Omega,\mathcal{F}_{0});H)$, we have that $(Y^{\varepsilon},\tilde{W})$ is a weak solution in the sense of [rock08, Def. 1.4] to

### v5

@@ 3.2. Local well-posedness
From now on, we let $\psi\in L^{2}(0,T;U)$ be arbitrary but fixed. Using Corollary 3.5, we will prove local well-posedness of the actual skeleton equation (2.8). Local well-posedness is established in Theorem 3.7. Its proof and preparatory lemma’s are analogous to [undefu, §18.2], which was inspired by [undefak] and [undefaj].
The skeleton equation does not fit in the setting of [undefu], [undefak] or [undefaj], for the reason that we only have $L^{1}$-( instead of $L^{2}$-)integrability of the term $B(\cdot,u^{\psi}(\cdot))\psi(\cdot)$ in (2.8). Besides that, our maximal regularity space $\mathrm{MR}(0,T)=C([0,T];H)\cap L^{2}(0,T;V)$ is different.
@@
[…]
@@
[…]
@@
[…]
@@
    $\leq C_{n,T}\Big(\int_{0}^{T}\|u(s)-v(s)\|_{H}^{2}\|w(s)\|_{V}^{2}\,\mathrm{d}s\Big)^{\frac{1}{2}}\leq C_{n,T}\|u-v\|_{C([0,T];H)}\|w\|_{L^{2}(0,T;V)}.$
@@
In (ii)-(v), note that strong measurability of $A_{0}(\cdot,u(\cdot))w(\cdot)$, $F(\cdot,u(\cdot))$, $B_{0}(\cdot,u(\cdot))w(\cdot)$ and $G(\cdot,u(\cdot))$ holds, as was mentioned in Remark 3.6. Moreover, by symmetry in Assumption 2.2, $B_{0}$ and $G$ can be estimated in the same way as $A_{0}$ and $F$. We provide the estimates for the latter.
Assumption 2.2(3) immediately yields (ii) and (iii).
@@
For (v): the following estimates can be found in the proof of [undefe, Prop. 4.5]. By Assumption 2.2(3) we have pointwise in $t\in[0,T]$:
@@
where $M_{\sigma,\beta,\rho}>0$ is a constant depending only on $\sigma$, $\beta$ and $\rho$ and we let $0^{0}=1$. In the above we used that $a\coloneqq\frac{(2\beta-1)\rho}{2-2\beta}\in[0,1]$, hence $x^{a}\leq 1+x$ for $x\geq 0$. For $j\in\{1,\ldots,m_{F}\}$, application of (3.2) gives pointwise in $t\in[0,T]$:
$\Big(1+\|$ $u\|_{\beta_{j}}^{\rho_{j}}+\|v\|_{\beta_{j}}^{\rho_{j}}\Big)\|u-v\|_{\beta_{j}}$
@@
[…]
@@
[…]
@@
The next lemma is analogous to [undefu, Lem. 18.2.10].
@@
where $\alpha_{T_{1}}(\tilde{T}),\beta_{T_{1},\sigma}(\tilde{T},r)\downarrow 0$ as $\tilde{T},r\downarrow 0$ $\alpha_{T_{1}}(\tilde{T}) and$ $\beta_{T_{1},\sigma}(\tilde{T},r)$ are independent of $v_{0}$ and $v$.
@@
[…]
@@
[…]
@@
Before we prove Theorem 3.7, we need one more lemma, a modification of [undefu, Lemma 18.2.12].
@@
[…]
@@
[…]
@@
$),\displaystyle\leq c_{T_{1}}\|v-w\|_{\mathrm{MR}(0,\tilde{T})}(r+\alpha(\tilde{T}$)),
@@
We are now ready to prove Theorem 3.7. The proof is adapted from [undefu, Th. 18.2.6].
@@
[…]
@@
with $\alpha_{T_{1}}(\tilde{T}),\beta_{T_{1},\sigma}(\tilde{T},r)\downarrow 0$ as $\tilde{T},r\downarrow 0$. Recall that $K_{T_{1}}$ from (3.8) only depends on $_{1}, TT$, $\|u_{0}\|_{H}$ and $\psi$, not on $v_{0}$ or $v$. Fixing first $\sigma\coloneqq(4K_{T_{1}})^{-1}$, we find
@@
[…]
@@
and note that $\delta\in(0,\min\{T_{0},\tilde{T}-s\}]$ since $v(0+s)=\tilde{v}(0+s)=w_{0}=z_{w_{0}}(0)$. In particular, $\delta\in(0,T_{0}]$ and $v(\cdot+s),\tilde{v}(\cdot+s)\in Z_{r_{0},T}(w_{0})$ by definition of $\delta$. Uniqueness of solutions in $Z_{r_{0},\delta}(w_{0})$ implies that $v(\cdot+s)=\tilde{v}(\cdot+s)$ on $0,\delta]$. Therefore $v=\tilde{v}$ on $0,s+\delta]$, contradicting the definition of $s$. We conclude that $v=\tilde{v}$. ∎
@@
[…]
$\geq\theta_{n,T}\|v\|_{V}^{2}-({M}_{n,T}+C_{\sigma}\|\psi(t)\|_{U}^{2})\|v\|_{H}^{2}-\sigma C_{n,T}^{2}\|v\|_{V}^{2}.$
Putting $\sigma\coloneqq\theta_{n,T}(2C_{n,T}^{2})^{-1}$, the required coercivity (3.3) for $\bar{A}$ follows.
@@ 3.3. Global well-posedness
Similar to [undefai, Chap. 5] and [undefu, §18.2], we will extend Theorem 3.7 to a global well-posedness result by means of maximal solutions and a blow-up criterion.
@@
Note that maximal solutions are unique by definition. The proof of the next proposition is adapted from [undefu, Th. 18.2.14, Th. 18.2.15] and [undefai].
@@
satisfies $\bar{u}\in\mathrm{MR}(0,T_{*}+{\delta})$ and $\bar{u}$ is a strong solution to (3.10) on $0,T_{*}+{\delta}]$, contradicting maximality of $(u_{*},T_{*})$. ∎
Using the blow-up criterion, we finally prove global well-posedness for the skeleton equation. Besides Assumption 2.2, we now also assume the coercivity condition (2.7) for the pair $(A,B)$. This condition has not been used so far, but it is also needed for the global well-posedness result for the stochastic evolution equation [undefe, Th. 3.5], see Theorem 2.3.
@@
[…]
@@
where $K\colon[0,T_{*}]\to{\mathbb{R}}_{+}$ is increasing. Applying Lemma 3.8(iv) with $n=K(T_{*})<\infty$ we find that $F(u_{*})\in L^{2}(0,t;V^{*})$ for all $t\in(0,T_{*})$ and $L\coloneqq\sup_{t\in[0,T_{*})}\|F(u_{*})\|_{L^{2}(0,t;V^{*})}<\infty$. Thus, by the Monotone Convergence Theorem, $\|F(u_{*})\|_{L^{2}(0,T_{*};V^{*})}\leq L<\infty$. Similarly, $G(u_{*})\in L^{2}(0,T_{*};\mathcal{L}_{2}(U,H))$. Now we apply Corollary 3.5 with $T\coloneqq T^{*}$, $w\coloneqq u_{*}\in C([0,T_{*});H)\subset L^{\infty}(0,T;H)$ (extend by $w(T)\coloneqq u_{*}(0)$ on the Lebesgue null set $\{T\}$), $n\coloneqq K(T_{*})$ and $\bar{f}\coloneqq F(u_{*})+f\in L^{2}(0,T_{*};V^{*})$, $\bar{g}\coloneqq G(u_{*})+g\in L^{2}(0,T_{*};\mathcal{L}_{2}(U,H))$. Corollary 3.5 gives existence of a strong solution $\bar{u}\in\mathrm{MR}(0,T_{*})$ to (2.8) on $0,T_{*}]$. By uniqueness of the maximal solution, it follows that $u_{*}|_{[0,T]}=\bar{u}|_{[0,T]}$ for all $T\in[0,T_{*})$. Hence $\lim_{t\uparrow T_{*}}u_{*}(t)=\lim_{t\uparrow T_{*}}\bar{u}(t)=\bar{u}(T_{*})\in H$, contradicting Proposition 3.15.
@@ 4.1. Weak convergence approach
    WfWf is normally distributed with mean zero and $\mathbb{E}[WfWg]=\langle f,g\rangle_{L^{2}({\mathbb{R}}_{+};U)}$,
@@
[…]
@@
[…]
@@
where $\mathcal{N}$ is the collection of all $(\Omega,\mathcal{F},\mathbb{P})$-null sets. Then $(\mathcal{H}_{t})_{t\geq 0}$ is a complete, right-continuous filtration on $(\Omega,\bar{\mathcal{F}},\bar{\mathbb{P}})$. Moreover, one can show that $W$ is a $U$-cylindrical Brownian motion with respect to $(\Omega,\bar{\mathcal{F}},\bar{\mathbb{P}},(\mathcal{H}_{t})_{t\geq 0})$ and with respect to $(\Omega,{\mathcal{F}},{\mathbb{P}},(\mathcal{F}^{0}_{t})_{t\geq 0})$. Let $Y^{\varepsilon}_{0}$ and $\bar{Y}^{\varepsilon}$ be the unique strong solution to (4.1) on $(\Omega,{\mathcal{F}},{\mathbb{P}},(\mathcal{F}^{0}_{t})_{t\geq 0})$ and $(\Omega,\bar{\mathcal{F}},\bar{\mathbb{P}},(\mathcal{H}_{t})_{t\geq 0})$, respectively. Since $\mathcal{F}^{0}_{t}\subset\mathcal{F}_{t}\cap\mathcal{H}_{t}$, $Y^{\varepsilon}_{0}$ is also a strong solution to (4.1) on $(\Omega,{\mathcal{F}},{\mathbb{P}},(\mathcal{F}_{t})_{t\geq 0})$ and on $(\Omega,\bar{\mathcal{F}},\bar{\mathbb{P}},(\mathcal{H}_{t})_{t\geq 0})$. Pathwise uniqueness gives ${Y}^{\varepsilon}=Y_{0}^{\varepsilon}=\bar{Y}^{\varepsilon}$ $\mathbb{P}$-a.s. Now trivially from Definition 2.4, if we prove the LDP for $(\bar{Y}^{\varepsilon})$, then the LDP carries over to $(Y^{\varepsilon})$.
@@
By [undefab, Prop. 2.5.2], $\tilde{W}_{1}$ is a $Q_{1}$-Wiener process on $U_{1}$, with $Q_{1}\coloneqq JJ^{*}$. In what follows, $\tilde{W}_{1}$ denotes this $Q_{1}$-Wiener process defined by (4.2). We note that the paths of $\tilde{W}_{1}$ are in $C([0,T];U_{1})$.
@@
$\mathcal{A}\coloneqq\{\Psi\colon[0,T]\times\Omega\to U:\Psi\text{ is an }(\mathcal{F}_{t})\text{-predictable process},\|\Psi\|_{L^{2}(0,T;U)}<\infty\;\mathbb{P}\text{-a.s.}\}$
and for $K>0$,
$S_{K}\coloneqq\{\psi\in L^{2}(0,T;U):\|\psi\|_{L^{2}(0,T;U)}\leq K\},\qquad\mathcal{A}_{K}\coloneqq\{\Psi\in\mathcal{A}:\Psi\in S_{K}\;\mathbb{P}\text{-a.s.}\}.$
The next theorem gives sufficient conditions for the LDP and is known as the weak convergence approach, which originates from [undefh, Th. 4.4]. In [undefac], a useful adaptation was proved. The following version is immediately derived from [undefac, Th. 3.2]. We will use it to prove Theorem 2.6.
Theorem 4.5.
@@
[…]
@@
Conditions (ii) and (iii) imply the conditions of the original weak convergence approach of [undefh]. For the latter, instead of (ii) and (iii), one would require
@@
Here, [undefh, Th. 4.4] is applied with $Q_{1}$-Wiener process $\tilde{W}_{1}$, $H\coloneqq U_{1}$, $H_{0}\coloneqq Q_{1}^{\frac{1}{2}}(U_{1})$ and one uses that $Q_{1}^{\frac{1}{2}}(U_{1})=J(U)=U$ as a subspace of $U_{1}$, see [undefab, Prop. 2.5.2] (with $Q\coloneqq I$, $U_{0}\coloneqq I^{\frac{1}{2}}(U)=U$).
@@
We will verify that all conditions in Theorem 4.5 are satisfied for $Y^{\varepsilon}$ defined as the strong solution to (4.1). Condition (i) follows from the Yamada-Watanabe theorem in [undefam]. The details are given in Lemma 4.6, as well as a preparation for the proof of condition (iii).
@@
[…]